OpenStudy (anonymous):

Find the supplementary angle of 79°

OpenStudy (anonymous):

It's 180 - 79 = 101

OpenStudy (anonymous):

in degrees, ofcourse :)

OpenStudy (anonymous):

supplement means two angles equal 180 degrees,, a+b=180 . if one angle is 79,,,79 +b = 180 b= 101 101 +79 =180 i dont no if what i have is good

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Any time :)